OpenStudy (anonymous):

log (x^4) - log (x^3)

OpenStudy (anonymous):

remember that if you have the same base, then \[log (a) - log(b)=log(\frac{a}{b})\]

OpenStudy (lone001):

log (x^4) - log (x^3)=log((x^4)/(x^3))=log(x)